OpenStudy (anonymous):

Trevor's dad decided to get him a cell phone for his birthday. Cell Plus has a plan that costs $24.00 per month plus an additional $0.06 per minute. Cell Best has a plan that costs $30.00 per month plus an additional $0.05 per minute. How many minutes can Trevor talk and both cell plans cost the same amount? 600 60 300 120

OpenStudy (nubeer):

minutes =x 24 + 0.06x Plan A 30+0.05x Plan B so both cost same... 24+0.06x = 30 +0.05x solve for x

OpenStudy (goformit100):

120

OpenStudy (anonymous):

explain

OpenStudy (mathstudent55):

Set up an expression for each plan, then set the expressions equal. Let m = number of minutes. Cell Plus: Fixed $24 + $0.06/minute 24 + 0.06m Cell Best: Fixed $30 + $0.05/minute 30 + 0.05m Solve the equation for m: 24 + 0.06m = 30 + 0.05m

OpenStudy (anonymous):

and it would all add up to 300 right ?

OpenStudy (anonymous):

the answer was 300. thanks for the help -_-

OpenStudy (mathstudent55):

24 + 0.06m = 30 + 0.05m - 0.05m - 0.05m ----------------------- 24 + 0.01m = 30 -24 -24 ------------------ 0.01m = 6 Divide both sides by 0.01 m = 600 It takes 600 minutes for both plans to cost the same.